​Lease-versus-purchase decision  Personal Finance Problem   Joanna Browne is considering either leasing or purchasing a new Chrysler Sebring convertible that has a​ manufacturer's suggested retail price​ (MSRP) of $33,500. The dealership offers a 3​-year lease that requires a capital payment of $3,374 ($3,074 down payment​ +$300 security​ deposit) and monthly payments of $495.

Purchasing requires a $2,620 down​ payment, sales tax of 6.9% ($2,312​), and 36 monthly payments of $922.

Joanna estimates the value of the car will be $17,000 at the end of 3 years. She can earn 4.6% annual interest on her savings and is subject to a 6.9% sales tax on purchases. Make a reasonable recommendation to Joanna using a​ lease-versus-purchase analysis​ that, for​ simplicity, ignores the time value of money.

a.  Calculate the total cost of leasing.

b.  Calculate the total cost of purchasing.

c.  Which should Joanna​ do?
Answer #1

a.  Calculate the total cost of leasing.

The total cost of leasing = Downpayment + Monthly lease payments x Lease perod in months + Opportunity cost on the downpayment over the lease eperiod = 3,374 + 495 x 3 years x 12 months + 3,374 x annual interest rate x 3 years = 3,374 + 495 x 36 + 3,374 x 4.6% x 3 = $  21,660

b.  Calculate the total cost of purchasing.

The total cost of purchasing = Downpayment + Sales tax payment + Monthly payments x Time in months + Opportunity cost on the downpayment over 3 years = 2,620 + 2,312 + 922 x 3 years x 12 months + 2,620 x annual interest rate x 3 years = 2,620 + 2,312 + 922 x 36 + 2,620 x 4.6% x 3 = $ 38,485

c.  Which should Joanna​ do?

The total cost of leasing < the total cost of purchasing. Hence Joanna should lease the car.